# Representational Alignment Across Model Layers and Brain Regions with Hierarchical Optimal Transport

### v1

@@ 4 Discussion
Hierarchical OT provides a principled framework for aligning two networks with arbitrary depths. By softly coupling all layer objectives, HOT (i) allows a neuron in one layer of a source network to align with a soft combination of units distributed across multiple layers of a target network, (ii) enforces global consistency that mitigates overfitting to noise in any single layer, and (iii) produces a single network-level—rather than layer-level—alignment score. In our experiments, HOT achieves higher alignment scores than greedy pairwise matches and yields intuitive transport plans that reveal hierarchical correspondences both in artificial networks and in the human brain.
@@
Despite these strengths, several limitations remain. First, HOT is computationally demanding: solving the inner OT scales as for neurons in a pair of layers, and the overall procedure is for networks with layers. This makes scaling to very wide or deep models challenging without further algorithmic improvements. Second, our evaluation was limited to a subset of models and brain datasets; follow-up analyses on more models and diverse neural data will be critical to assess generality. Finally, while HOT provides a descriptive alignment, it does not by itself explain why certain features converge across systems, or which computational principles drive these correspondences.
@@
Several extensions are natural. One avenue is to add a third hierarchical level to capture representational dynamics over the course of training, enabling alignment not only across neurons and layers but also across developmental or learning trajectories. Applications could include aligning models across different stages of training, or comparing biological learning to gradient descent. Another is to incorporate priors on the transport plan (e.g. smoothness) to guide alignment toward more interpretable solutions. Ultimately, as both biological and artificial networks grow in scale and complexity, methods like HOT that respect global structure while revealing network correspondences will be essential for understanding the universal principles governing intelligent systems.
@@
Appendix A Comparing LLM-LLM Representations
Appendix B Representation Similarity between Vision Cortex Response
| Model 1 | Model 2 | HOT Metric | Random (Perm-P) | Single-Best OT | Pairwise Best OT |
| Subject A | Subject B | 0.306 0.007 | 0.258 0.008 | 0.306 0.007 | 0.308 0.007 |
| Subject A | Subject C | 0.227 0.008 | 0.184 0.008 | 0.227 0.008 | 0.232 0.009 |
| Subject A | Subject D | 0.248 0.014 | 0.212 0.016 | 0.248 0.014 | 0.253 0.014 |
| Subject B | Subject C | 0.237 0.008 | 0.199 0.006 | 0.237 0.008 | 0.237 0.009 |
| Subject C | Subject D | 0.235 0.010 | 0.196 0.009 | 0.235 0.010 | 0.239 0.008 |
| Subject B | Subject D | 0.248 0.013 | 0.217 0.011 | 0.248 0.013 | 0.254 0.012 |
B.1 Transport plans for OT based mapping pairs
B.2 Transport plans for Linear Predictivity based mapping pairs
Appendix C Representation Similarity between Vision Models
| Model 1 | Model 2 | HOT Metric | Random (Perm-P) | Single-Best OT | Pairwise Best OT |
@@ Appendix C Representation Similarity between Vision Models
| Model 1 | Model 2 | HOT + R | Single-Best + R | Pairwise Best + R |
@@
Appendix D Representation Similarity between Vision Models and Vision Cortex
Experimental Setup. We assess representational similarity between human visual cortex and vision transformers by comparing fMRI responses with model activations elicited by the same set of stimuli. Specifically, we use responses from the Natural Scenes Dataset ([Allen et al., 2022]), focusing on 1,000 shared images viewed by participants in the fMRI experiment. The same images are presented to pretrained vision transformers, and layer-wise representations are extracted by averaging patch embeddings across each input. Following the approach used in the cortex–cortex analysis, we treat distinct visual areas (V1–V4) as “layers” and individual voxels as “neurons.” We then compute both HOT and its rotation-augmented variant (HOT+R) to align cortical responses with model representations. This design enables a direct comparison of hierarchical organization across biological and artificial systems under matched visual input.
| Model 1 | Model 2 | HOT Metric | Pairwise Best OT | HOT + R | Pairwise Best + R |
@@ Appendix D Representation Similarity between Vision Models and Vision Cortex
Results. Table D.1 reports reconstruction scores for HOT, HOT+R, and corresponding baselines averaged across the four subjects. Among all methods, HOT+R achieves the highest reconstruction score, indicating that incorporating rotation is critical for capturing shared structure between cortical and model representations. Subject-specific results (Tables D.2 and D.3) further confirm this trend, with HOT+R consistently outperforming both vanilla HOT and pairwise baselines. Transport plans visualized in Figures D.1– D.16 reveal partial but inconsistent layer-wise correspondences: in some cases, early cortical regions align with early model layers and higher regions map to deeper layers, while in others the mappings appear noisier.
| Model 1 | Model 2 | HOT Metric | Random (Perm-P) | Single-Best OT | Pairwise Best OT |
@@
| Model 1 | Model 2 | HOT + R | Single-Best + R | Pairwise Best + R |
@@
Appendix E Use of Large Language Models

### v2

@@ 4 Discussion
Multi-Level OT provides a principled framework for aligning two networks with arbitrary depths. By softly coupling all layer objectives, MOT (i) allows a neuron in one layer of a source network to align with a soft combination of units distributed across multiple layers of a target network, (ii) enforces global consistency that mitigates overfitting to noise in any single layer, and (iii) produces a single network-level—rather than layer-level—alignment score. In our experiments, MOT achieves higher alignment scores than greedy pairwise matches and yields intuitive transport plans that reveal hierarchical correspondences both in artificial networks and in the human brain.
@@
Despite these strengths, several limitations remain. First, MOT is computationally demanding: solving the inner OT scales as for neurons in a pair of layers, and the overall procedure is for networks with layers. This makes scaling to very wide or deep models challenging without further algorithmic improvements. Second, our evaluation was limited to a subset of models and brain datasets; follow-up analyses on more models and diverse neural data will be critical to assess generality. Finally, our focus here is on developing a principled method for quantifying representational alignment, rather than explaining why different systems converge on similar representations. The latter is a separate and substantially harder theoretical question. Recent ideas, such as the contravariance principle (which frames convergence as a consequence of shared task or ecological constraints) (Cao and Yamins, 2024) and the Platonic representation hypothesis (which proposes that training on rich naturalistic data drives different systems toward similar underlying structure in the world) (Huh et al., 2024), offer intriguing conceptual directions but remain challenging to evaluate empirically. By providing a rigorous way to characterize alignment, MOT offers the measurement foundation needed for such explanatory hypotheses to be tested in future work.
@@
Several extensions are natural. One is to study training-run level alignment in more depth, examining how model architecture, training data order, learning rate, and other optimization choices shape the resulting alignment plan. Another is to incorporate priors on the transport plan (e.g. smoothness) to guide alignment toward more interpretable solutions. Ultimately, as both biological and artificial networks grow in scale and complexity, methods like MOT that respect global structure while revealing network correspondences will be essential for understanding the universal principles governing intelligent systems.
@@
Appendix A Optimal Transport for Representational Alignment
Optimal Transport (OT) provides a principled way to compare two neural populations by framing each representation as a probability distribution over its tuning functions and measuring the minimal “effort” required to transform one representational distribution into another. Concretely, if and denote two sets of tuning curves across stimuli, we represent each population as a uniform mixture of Dirac masses over its neurons:
The cost function quantifies the dissimilarity between the tuning functions of neuron in and neuron in ; depending on the application, may be squared Euclidean distance, correlation distance, or any appropriate tuning-based dissimilarity. The 2-Wasserstein or the soft-matching (Khosla and Williams [2024]) distance between and is
where ranges over the transportation polytope Any is a doubly stochastic coupling: it assigns a non-negative “mass” to transporting representational weight from neuron in to neuron in , with uniform row and column marginals. Unlike rotation-invariant metrics such as RSA or CKA, OT preserves single-neuron tuning structure and yields explicit, interpretable neuron-to-neuron correspondences.
MOT builds directly on this OT foundation by elevating the coupling to the level of layers, simultaneously inferring soft layer-to-layer couplings and per-neuron transport plans to produce a globally balanced alignment across networks of arbitrary depths (Figure 1).
Appendix B Comparing LLM-LLM Representations
Appendix C Representation Similarity between Vision Cortex Response
| Model 1 | Model 2 | MOT Metric | Random (Perm-P) | Single-Best OT | Pairwise Best OT |
| Subject A | Subject B | 0.244 0.005 | 0.135 0.015 | 0.244 0.005 | 0.245 0.004 |
| Subject A | Subject C | 0.199 0.003 | 0.110 0.013 | 0.199 0.003 | 0.202 0.003 |
| Subject A | Subject D | 0.198 0.008 | 0.109 0.016 | 0.198 0.008 | 0.201 0.007 |
| Subject B | Subject C | 0.212 0.007 | 0.126 0.014 | 0.212 0.007 | 0.212 0.007 |
| Subject C | Subject D | 0.197 0.007 | 0.112 0.016 | 0.197 0.007 | 0.199 0.006 |
| Subject B | Subject D | 0.201 0.008 | 0.121 0.014 | 0.201 0.008 | 0.204 0.007 |
C.1 Transport plans for OT based mapping pairs
Appendix D Representation Similarity between Vision Models
| Model 1 | Model 2 | MOT Metric | Random (Perm-P) | Single-Best OT | Pairwise Best OT |
@@ Appendix D Representation Similarity between Vision Models
| Model 1 | Model 2 | MOT + R | Single-Best + R | Pairwise Best + R |
@@
Appendix E Representation Similarity between Vision Models and Vision Cortex
Experimental Setup. We assess representational similarity between human visual cortex and vision transformers by comparing fMRI responses with model activations elicited by the same set of stimuli. Specifically, we use responses from the Natural Scenes Dataset ([Allen et al., 2022]), focusing on 1,000 shared images viewed by participants in the fMRI experiment. The same images are presented to pretrained vision transformers, and layer-wise representations are extracted by averaging patch embeddings across each input. Following the approach used in the cortex–cortex analysis, we treat distinct visual areas (V1–V4) as “layers” and individual voxels as “neurons.” We then compute both MOT and its rotation-augmented variant (MOT+R) to align cortical responses with model representations. This design enables a direct comparison of hierarchical organization across biological and artificial systems under matched visual input.
| Model 1 | Model 2 | MOT Metric | Pairwise Best OT | MOT + R | Pairwise Best + R |
@@ Appendix E Representation Similarity between Vision Models and Vision Cortex
Results. Table E.1 reports reconstruction scores for MOT, MOT+R, and corresponding baselines averaged across the four subjects. Among all methods, MOT+R achieves the highest reconstruction score, indicating that incorporating rotation is critical for capturing shared structure between cortical and model representations. Subject-specific results (Tables E.2 and E.3) further confirm this trend, with MOT+R consistently outperforming both vanilla MOT and pairwise baselines. Transport plans visualized in Figures E.1– E.16 reveal partial but inconsistent layer-wise correspondences: in some cases, early cortical regions align with early model layers and higher regions map to deeper layers, while in others the mappings appear noisier.
| Model 1 | Model 2 | MOT Metric | Random (Perm-P) | Single-Best OT | Pairwise Best OT |
@@
| Model 1 | Model 2 | MOT + R | Single-Best + R | Pairwise Best + R |
@@
Appendix F Global Alignment
In this section, we report reconstruction scores on a held-out validation dataset for the vision-cortex alignment study under a global alignment formulation. In this setting, all layers from both networks are flattened so that all neurons are treated as belonging to a single layer. We compare the resulting reconstruction scores to those obtained with our MOT reconstruction scores.
Global alignment scores are computed using two alignment metrics: optimal transport and linear predictivity. Reconstruction performance is then evaluated on held-out validation data. The results for this experiment are presented in Table F.1.
| Model 1 | Model 2 | Global OT | Global Linear Predictivity | MOT |
| Subject A | Subject B | 0.184 | 0.305 | 0.244 |
| Subject A | Subject C | 0.177 | 0.295 | 0.199 |
| Subject A | Subject D | 0.168 | 0.289 | 0.198 |
| Subject B | Subject C | 0.183 | 0.306 | 0.212 |
| Subject C | Subject D | 0.158 | 0.306 | 0.197 |
| Subject B | Subject D | 0.163 | 0.297 | 0.201 |
Appendix G Quantifying Hierarchy
Given a layer-to-layer transport plan , we define a scalar “hierarchy” score that quantifies how strongly the transport mass is concentrated along a diagonal correspondence between source and target layers.
We first embed the layer indices into the unit interval via
so that the ideal diagonal correspondence between layers lies along the line . For each entry we define its normalized distance to this diagonal as
We interpret the magnitudes of the transport plan as a discrete mass distribution,
and compute the average distance of this mass from the diagonal,
Finally, we define the hierarchy score as
By construction and it depends only on the relative placement of transport mass with respect to the diagonal. Table G.1 reports for each transport plan as a summary measure of its hierarchy.
| Model 1 | Model 2 | MOT | Pairwise OT | MOT+R | Pairwise Best + R |
| DINOv2 Small | ViT-MAE Large | 0.90 | |||
| ViT-MAE Base | ViT-MAE Large | 0.97 | |||
| DINOv2 Small | ViT-MAE Base | 1.00 | |||
| ViT-MAE Base | DINOv2 Large | 0.96 | |||
| DINOv2 Small | DINOv2 Large | 0.98 | |||
| ViT-MAE Huge | DINOv2 Giant | 0.86 | |||
| ViT-MAE Base | DINOv2 Giant | 0.91 | |||
| DINOv2 Small | DINOv2 Giant | 0.97 | |||
| ViT-MAE Base | ViT-MAE Huge | 0.83 | |||
| DINOv2 Small | ViT-MAE Huge | 0.88 | |||
| Llama-3.2 1B | Llama-3.2 3B | 0.96 | |||
| Qwen-2.5 0.5B | Llama-3.2 3B | 0.86 | |||
| Llama-3.2 3B | Qwen-2.5 3B | 0.83 | |||
| Llama-3.2 1B | Qwen-2.5 3B | 0.85 | |||
| Qwen-2.5 0.5B | Llama-3.2 1B | 0.86 | |||
| Qwen-2.5 0.5B | Qwen-2.5 3B | 0.93 | |||
| Subject B | Subject D | 1.00 | |||
| Subject C | Subject D | 1.00 | |||
| Subject B | Subject C | 1.00 | |||
| Subject A | Subject B | 1.00 | |||
| Subject A | Subject C | 1.00 | |||
| Subject A | Subject D | 1.00 |
We observe that, across all transport plans, the hierarchy metric is highest for plans generated by MOT (or MOT+R), compared with those produced by pairwise best (or pairwise rotational) methods. This indicates that MOT and MOT+R based approaches uncover hierarchical correspondences that pairwise methods don’t reveal.
Appendix H Transport plans for Linear Predictivity based mapping pairs
| Model 1 | Model 2 | Pairwise LP Reconstruction score | MOT Metric |
| Llama-3.2 1B | Llama-3.2 3B | ||
| Qwen-2.5 0.5B | Llama-3.2 3B | ||
| Llama-3.2 3B | Qwen-2.5 3B | ||
| Llama-3.2 1B | Qwen-2.5 3B | ||
| Qwen-2.5 0.5B | Llama-3.2 1B | ||
| Qwen-2.5 0.5B | Qwen-2.5 3B | ||
| Subject B | Subject D | ||
| Subject C | Subject D | ||
| Subject B | Subject C | ||
| Subject A | Subject B | ||
| Subject A | Subject C | ||
| Subject A | Subject D |
Appendix I Transport plans for RSA-based mapping pairs
Appendix J Robustness of metrics to sub-sampling
We consider a pair of LLMs (Qwen-2.5 0.5B and Llama-3.2 1B) and systematically sub-sample neurons from the larger model (Llama3.2 1B) to study how this affects reconstruction of the smaller model’s activations. For each fraction of neurons kept, we randomly select that fraction of neurons from Llama-3.2 1B, re-fit the transport plans between the two models, and evaluate the reconstruction score on held-out test data (a 20% split, as before), repeating this procedure over five different seeds. Figure J.1 shows the resulting reconstruction scores for MOT and MOT+R (with their corresponding pairwise baselines).
We observe that reconstruction scores for both MOT/MOT+R and the corresponding baselines increase with the fraction of neurons kept. Importantly, sub-sampling does not cause MOT to collapse; its performance degrades smoothly. For MOT+R, keeping only 10% of neurons yields a reconstruction score of nearly 0.5, indicating substantial robustness to aggressive sub-sampling. In the upper range of fraction of neurons kept, MOT+R achieves substantially higher reconstruction scores than both its rotation-aware baseline and the corresponding pairwise OT and MOT scores. These results suggest that neuron sub-sampling is a practical way to trade off computation and accuracy, and that the metrics remain informative even under significant sub-sampling.
Appendix K Use of Large Language Models